# Generation-Time vs. Post-hoc Citation: A Holistic Evaluation of LLM Attribution

### v1

@@ Abstract
Trustworthy Large Language Models (LLMs) must cite human-verifiable sources in high-stakes domains such as healthcare, law, scientific research, where even small errors can have severe consequences. Practitioners and researchers face a choice: let models generate citations during decoding, or let models draft answers first and then attach appropriate citations. To clarify this choice, we introduce two paradigms: Generation-Time Citation (G-Cite), which produces the answer and citations in one pass, and Post-hoc Citation (P-Cite), which adds or verifies citations after drafting. We conduct a comprehensive evaluation from zero-shot to advanced retrieval-augmented methods across four popular attribution datasets, and provide evidence-based recommendations that weigh trade-offs across use cases. Our results show a consistent trade-off between coverage and citation correctness, with retrieval as the main driver of attribution quality in both paradigms. P-Cite methods achieve high coverage with competitive correctness and moderate latency, whereas G-Cite methods prioritize precision at the cost of coverage and speed. We recommend a retrieval-centric, P-Cite-first approach for high-stakes applications, reserving G-Cite for precision-critical settings such as strict claim verification. Our codes and human evaluation results are available at https://anonymous.4open.science/r/Citation_Paradigms-BBB5/
@@ 1 Introduction
Generation-Time Citation (G-Cite) creates the text and citation markers together in a single step. Post-hoc Citation (P-Cite) works differently; it first creates a draft, then adds or checks citations in a separate step. The key difference between these approaches is timing of citations. Beside timings, these approaches differ in how they work technically. G-Cite makes citation choices during the normal left-to-right text generation process. It decides locally based on what has been written so far and any retrieved evidence. In contrast, P-Cite separates citation from text generation entirely. It runs a second pass to examines the complete draft and available evidence to add or verify citations throughout the entire text.
@@
Table 1: Details of Datasets used in the Experiments. FEVER focus on factual verification of claims in health, law and other domains. REASONS is focused on scientific research, and LongCite/ALCE are open domain QA with long and short context respectively. “Doc”, “Sent‘”, “Claim” denotes if the citation is at document-level, sentence-level, or Claim-level.
As of now there is no principle and systematic study to compare these two paradigm to determine its limitation and capabilities. Each method is designed differently and evaluated using a different evaluation metric on different datasets. These gap makes it difficult for researchers and practitioners to choose the right method or design better attribution systems. To address this problem, our paper provides a rigorous empirical comparison of both paradigm using common datasets and evaluation metric across diverse categories of methods.
@@
We benchmark the state-of-the-art methods from both paradigms. Because several existing datasets are tailored to a single paradigm, we adapt them to enable fair, cross-paradigm comparison. More details on the datasets are available in Table 1. We consider four types of methods: Zero-shot, Fine-tuned, Retrieval Augmented Generation (RAG) and two recent approaches from G-Cite and P-Cite as Advanced methods. Here zero-shot and RAG act as common baselines. For Advanced and Fine-tuned methods, within G-Cite, we use CoT Citation (Ji et al., 2024) and LongCite (8b) (Zhang et al., 2025) and within P-Cite, we evaluate CiteBART (Çelik and Tekir, 2025) and CEG (Li et al., 2024a). We assess performance using established quantitative attribution metrics: Citation Correctness, Precision, Recall, Coverage, and Latency. These metrics are widely used in prior work and enable consistent comparison across methods. We also conduct a human evaluation with n=100 instances per method-dataset pair, providing 80% statistical power to detect medium effect sizes (Cohen’s d $\geq$ 0.5) at $\alpha$ = 0.05 significance level, for more details, see section 3.
@@ 2.1 Generation-time Citation (G-Cite)
Datasets: In open domain short-context datasets, where the documents containing the citation to the correct answer are relatively short, include ASQA (Stelmakh et al., 2022), QAMPARI (Amouyal et al., 2023), and ALCE. For long-context datasets, where the length of the documents containing the citation is relatively large, there exist popular datasets like ELI5 (Fan et al., 2019) and LongBench-Cite.
@@ 2.2 Post-hoc Citation (P-Cite)
Methods: The second category, P-Cite, includes methods that generate citations for a pre-existing text. A prominent example is the REASONS benchmark (Saxena et al., 2025), which provides a dataset and Retrieval-Augmented Generation (RAG) baselines for sentence-level citation. Other methods also leverage RAG, using iterative approaches (Li et al., 2024a) or adding a verification step for each claim, as seen in RARR. Another recent approach that builds on top of RAG systems is CiteFix (Maheshwari et al., 2025), it uses a two step citation correction method to achieve better citation quality. Beyond standard RAG, approaches in this paradigm include fine-tuning with models like CiteBART (Çelik and Tekir, 2025) and using hierarchical attention mechanisms as in HAtten (Gu et al., 2022). Research in P-Cite also addresses related sub-tasks, such as recommending missed citations (Long et al., 2024) or determining if a sentence requires a citation at all (Batista et al., 2025b).
@@ 3 Experimentation
Methods. We evaluate G-Cite and P-Cite using a systematic categorization of four method types: (i) Zero-shot prompting, (ii) Fine-tuned models, (iii) Retrieval-augmented methods, and (iv) Advanced hybrid techniques. Zero-shot prompting was used across both the G-Cite and P-Cite paragraph, serving as a fundamental baselines without task-specific training. Implementation details and zero-shot prompts with examples are in subsection A.1. Fine-tuned models represent specialized architectures trained on domain-specific datasets such as LongCite-8B for citation generation and CiteBART for paper citation, both are trained using supervised learning on ALCE and arXiv datasets. Retrieval-augmented methods combine retrieval systems with generative models to enhance citation accuracy and coverage. We considered retrieval-based methods in both G-Cite and P-Cite. Advanced hybrid techniques integrate multiple methodological components: CoT Citation combines evidence retrieval with chain-of-thought prompting and includes an evidence-insurance step for comprehensive citation coverage, while Citation Evidence Generation (CEG) employs an iterative approach that retrieves relevant information before systematically attaching and verifying citations.
@@
Figure 2: Human Evaluation Results. We report Answer Correctness ($\uparrow$), and Citation Hallucination ($\downarrow$), values are averaged over all datasets and methods within each paradigm (G-Cite and P-Cite).P-Cite based methods tends to provide more correct answers with lesser hallucination.
@@
Metrics and Human Evaluation. We use standard citation metrics for evaluation. Specifically, we use five standard metrics: (i) Citation Precision to measure the fraction of correct citations among those produced; ii) Citation Recall to measure the fraction of ground-truth citations that are retrieved;(iii) Citation Correctness to measure the harmonic mean of precision and recall (Aly et al., 2024); (iv) Coverage to measure the proportion of ground-truth citations present in the generated response (Aly et al., 2024); (v) Latency to measure the average time (in seconds) taken by each method per dataset instance. We conducted a human assessment using two expert annotators from the university library with experience in AI-assisted citation verification. We evaluated 100 instances per method-dataset pair across two critical quality measures for each generated response, achieving $\kappa$ = 0.873 inter-annotator agreement. Answer Correctness is a strict metric which allow human evaluators to verify whether the provided evidence actually supports each claim made in the generated text. We assign a score of 1 when all claims are properly backed by their cited sources, and 0 when any claims lack adequate support. Citation Hallucination allows human evaluators to check whether each citation corresponds to a real source from the reference dataset. Humans assign a score of 1 when citations are fabricated or point to sources outside the ground truth collection, and 0 when all citations are legitimate and verifiable.
@@ 4 Results and Analysis
Our evaluation reveals distinct performance characteristics between the two citation paradigms across all datasets. P-Cite consistently achieve higher coverage while maintaining competitive citation correctness compared to G-Cite. For practitioners deploying LLMs in information-seeking applications, where users need comprehensive source attribution to verify claims across multiple documents, P-Cite methods provide a critical advantage by ensuring broader citation coverage without compromising accuracy. Further, as shown in Figure 2, the human evaluation reinforces these findings: averaged over datasets and methods, P-Cite shows higher answer correctness than G-Cite (78% vs. 69%) and lower citation hallucination (37% vs. 41%). See Appendix A.2 for detailed results.
@@
Finding 2: On the scientific REASONS dataset both paradigms achieve comparable correctness (P-Cite 26% vs. G-Cite 27%) and near-ceiling coverage (P-Cite 99% vs. G-Cite: 97%). On the FEVER dataset, G-Cite achieves the highest precision and correctness (94%) but limited citation coverage (27%), while P-Cite provides a more balanced profile with high coverage (74%) and strong correctness (75%). This shows that in scientific literature tasks, both P-cite and G-cite methods perform similarly well. However, for fact verification tasks, which are important for legal and policy-driven research, P-cite methods excel when you need comprehensive evidence from multiple sources, while G-cite methods work better when you need extremely precise validation of individual claims.
@@
Finding 4: Advanced methods built on retrieval foundations adjust the citation coverage and latency trade-off. P-Cite delivers high coverage and correctness with moderate latency costs, whereas G-Cite delivers better performance but substantially increases latency. Practitioners must balance operational efficiency with accuracy, where P-Cite offer a practical solution and G-Cite could be used for verification purposes. Fine-tuned models provide incremental improvements but cannot replace retrieval for maintaining content accuracy. Human evaluators, representing end users who rely on generated content for critical decisions, consistently rate P-Cite outputs as more accurate and trustworthy.

### v2

@@ Abstract
Trustworthy Large Language Models (LLMs) must cite human-verifiable sources in high-stakes domains such as healthcare, law, and scientific research, where even small errors can have severe consequences. Practitioners and researchers face a choice: let models generate citations during decoding, or let models draft answers first and then attach appropriate citations. To clarify this choice, we introduce two paradigms: Generation-Time Citation (G-Cite), which produces the answer and citations in one pass, and Post-hoc Citation (P-Cite), which adds or verifies citations after drafting. We conduct a comprehensive evaluation from zero-shot to advanced retrieval-augmented methods across four popular attribution datasets, and provide evidence-based recommendations that weigh trade-offs across use cases. Our results show a consistent trade-off between coverage and citation correctness, with retrieval as the main driver of attribution quality in both paradigms. P-Cite methods achieve high coverage with competitive correctness and moderate latency, whereas G-Cite methods prioritize precision at the cost of coverage and speed. We recommend a retrieval-centric, P-Cite-first approach for high-stakes applications, reserving G-Cite for precision-critical settings such as strict claim verification. Our codes and human evaluation results are available at https://anonymous.4open.science/r/Citation_Paradigms-BBB5/
@@ 1 Introduction
Generation-Time Citation (G-Cite) creates the text and citation markers together in a single step. Post-hoc Citation (P-Cite) works differently; it first creates a draft, then adds or checks citations in a separate step. The key difference between these approaches is the timing of citations. Besides timings, these approaches differ in how they work technically. G-Cite makes citation choices during the normal left-to-right text generation process. It decides locally based on what has been written so far and any retrieved evidence. In contrast, P-Cite separates citation from text generation entirely. It runs a second pass to examine the complete draft and available evidence, adding or verifying citations throughout the text.
@@
Table 1: Details of Datasets used in the Experiments. FEVER focuses on factual verification of claims in health, law, and other domains. REASONS is focused on scientific research, and LongCite/ALCE are open domain QA with long and short context, respectively. “Doc”, “Sent”, “Claim” denote if the citation is at document-level, sentence-level, or Claim-level.
As of now, there is no principal and systematic study to compare these two paradigms to determine their limitations and capabilities. Each method is designed differently and evaluated using a different evaluation metric on different datasets. These gap makes it difficult for researchers and practitioners to choose the right method or design better attribution systems. To address this problem, our paper provides a rigorous empirical comparison of both paradigms using standard datasets and evaluation metrics across diverse categories of methods.
@@
We benchmark the state-of-the-art methods from both paradigms. Because several existing datasets are tailored to a single paradigm, we adapt them to enable fair, cross-paradigm comparison. More details on the datasets are available in Table 1. We consider four types of methods: Zero-shot, Fine-tuned, Retrieval Augmented Generation (RAG), and two recent approaches from G-Cite and P-Cite as Advanced methods. Here, zero-shot and RAG act as common baselines. For Advanced and Fine-tuned methods, within G-Cite, we use CoT Citation (Ji et al., 2024) and LongCite (8b) (Zhang et al., 2025) and within P-Cite, we evaluate CiteBART (Çelik and Tekir, 2025) and Citation-Enhanced Generation (CEG) (Li et al., 2024a). We assess performance using established quantitative attribution metrics: Citation Correctness, Precision, Recall, Coverage, and Latency. These metrics are widely used in prior work and enable consistent comparison across methods. We also conduct a human evaluation with n=100 instances per method-dataset pair, providing 80% statistical power to detect medium effect sizes (Cohen’s d $\geq$ 0.5) at $\alpha$ = 0.05 significance level, for more details, see section 3.
@@ 2.1 Generation-time Citation (G-Cite)
Datasets: In open domain short-context datasets, where the documents containing the citation to the correct answer are relatively short, include ASQA (Stelmakh et al., 2022), QAMPARI (Amouyal et al., 2023), and ALCE. For long-context datasets, where the documents containing the citation are relatively long, there are popular datasets such as ELI5 (Fan et al., 2019) and LongBench-Cite.
@@ 2.2 Post-hoc Citation (P-Cite)
Methods: The second category, P-Cite, includes methods that generate citations for a pre-existing text. A prominent example is the REASONS benchmark (Saxena et al., 2025), which provides a dataset and Retrieval-Augmented Generation (RAG) baselines for sentence-level citation. Other methods also leverage RAG, using iterative approaches (Li et al., 2024a) or adding a verification step for each claim, as seen in RARR. Another recent approach that builds on top of RAG systems is CiteFix (Maheshwari et al., 2025), which uses a two-step citation correction method to achieve better citation quality. Beyond standard RAG, approaches in this paradigm include fine-tuning with models like CiteBART (Çelik and Tekir, 2025) and using hierarchical attention mechanisms as in HAtten (Gu et al., 2022). Research in P-Cite also addresses related sub-tasks, such as recommending missed citations (Long et al., 2024) or determining if a sentence requires a citation at all (Batista et al., 2025b).
@@ 3 Experimentation
Methods. We evaluate G-Cite and P-Cite using a systematic categorization of four method types: (i) Zero-shot prompting, (ii) Fine-tuned models, (iii) Retrieval-augmented methods, and (iv) Advanced hybrid techniques. Zero-shot prompting was used across both the G-Cite and P-Cite paragraphs, serving as a fundamental baseline without task-specific training or fine-tuning. Implementation details and zero-shot prompts with examples are in subsection A.1. Fine-tuned models represent specialized architectures trained on domain-specific datasets, such as LongCite-8B for citation generation and CiteBART for paper citations, both trained using supervised learning on ALCE and arXiv datasets. Retrieval-augmented methods combine retrieval systems with generative models to enhance citation accuracy and coverage. We considered retrieval-based methods in both G-Cite and P-Cite. Advanced hybrid techniques integrate multiple methodological components: CoT Citation combines evidence retrieval with chain-of-thought prompting and includes an evidence-insurance step for comprehensive citation coverage, while CEG employs an iterative approach that retrieves relevant information before systematically attaching and verifying citations.
@@
Figure 2: Human Evaluation Results. We report Answer Correctness ($\uparrow$), and Citation Hallucination ($\downarrow$), values are averaged over all datasets and methods within each paradigm (G-Cite and P-Cite).P-Cite based methods tend to provide more correct answers with lesser hallucination.
@@
Metrics and Human Evaluation. We use standard citation metrics for evaluation. Specifically, we use five standard metrics: (i) Citation Precision to measure the fraction of correct citations among those produced; ii) Citation Recall to measure the fraction of ground-truth citations that are retrieved;(iii) Citation Correctness to measure the harmonic mean of precision and recall (Aly et al., 2024); (iv) Coverage to measure the proportion of ground-truth citations present in the generated response (Aly et al., 2024); (v) Latency to measure the average time (in seconds) taken by each method per dataset instance. We conducted a human assessment using two expert annotators from the university library with experience in AI-assisted citation verification. We evaluated 100 instances per method-dataset pair across two critical quality measures for each generated response, achieving $\kappa$ = 0.873 inter-annotator agreement. Answer Correctness is a strict metric that allows human evaluators to verify whether the provided evidence actually supports each claim made in the generated text. We assign a score of 1 when all claims are properly backed by their cited sources, and 0 when any claims lack adequate support. Citation Hallucination allows human evaluators to check whether each citation corresponds to a real source from the reference dataset. Humans assign a score of 1 when citations are fabricated or point to sources outside the ground truth collection, and 0 when all citations are legitimate and verifiable.
@@ 4 Results and Analysis
Our evaluation reveals distinct performance characteristics between the two citation paradigms across all datasets. P-Cite consistently achieves higher coverage while maintaining competitive citation correctness compared to G-Cite. For practitioners deploying LLMs in information-seeking applications, where users need comprehensive source attribution to verify claims across multiple documents, P-Cite methods provide a critical advantage by ensuring broader citation coverage without compromising accuracy. Further, as shown in Figure 2, the human evaluation reinforces these findings: averaged over datasets and methods, P-Cite shows higher answer correctness than G-Cite (78% vs. 69%) and lower citation hallucination (37% vs. 41%). See Appendix A.2 for detailed results.
@@
Finding 2: On the scientific REASONS dataset, both paradigms achieve comparable correctness (P-Cite 26% vs. G-Cite 27%) and near-ceiling coverage (P-Cite 99% vs. G-Cite: 97%). On the FEVER dataset, G-Cite achieves the highest precision and correctness (94%) but limited citation coverage (27%), while P-Cite provides a more balanced profile with high coverage (74%) and strong correctness (75%). This shows that in scientific literature tasks, both P-cite and G-cite methods perform similarly well. However, for fact verification tasks, which are important for legal and policy-driven research, P-cite methods excel when you need comprehensive evidence from multiple sources, while G-cite methods work better when you need extremely precise validation of individual claims.
@@
Finding 4: Advanced methods built on retrieval foundations adjust the citation coverage and latency trade-off. P-Cite delivers high coverage and correctness with moderate latency costs, whereas G-Cite delivers better performance but substantially increases latency. Practitioners must balance operational efficiency with accuracy, where P-Cite offers a practical solution and G-Cite could be used for verification purposes. Fine-tuned models provide incremental improvements but cannot replace retrieval for maintaining content accuracy. Human evaluators, representing end users who rely on generated content for critical decisions, consistently rate P-Cite outputs as more accurate and trustworthy.
@@ 5 Conclusion
6 Acknowledgment
We thank the human annotators who contributed to this study for their valuable efforts. We also thank the anonymous reviewers for their constructive feedback and suggestions, which helped improve the quality of this work. We gratefully acknowledge support from the UMBC Faculty Start-up, Cybersecurity Leadership-Exploratory Grant and the USISTEF Award. The opinions, conclusions, and recommendations expressed here are solely those of the authors and do not necessarily reflect the views of USISTEF or UMBC.